# EEG-Based Consumer Behavior Prediction: An Exploration from Classical Machine Learning to Graph Neural Networks

### v1

@@ 1 Introduction
As purchasing decisions and consumer behaviour play a crucial role and have a significant impact on the success or failure of a business, a deeper understanding of this process has become particularly important for marketing [13], [14], [15], [16] and product development [17], [18]. Companies and organizations can optimize their marketing strategies [19], personalize advertisements for their consumers, and present a better user experience [20] by analyzing the way their consumers behave. Especially, in today’s competitive markets [21], the ability to predict consumers’ reactions to marketing campaigns and newly released products would make a competitive advantage and could cause a decrease in costs [22].
Many different methods are used to study and analyse brain activities that are related to the decision-making process and consumer behaviour. These methods include Functional magnetic resonance imaging (fMRI) [23], [24], Functional near-infrared spectroscopy (fNIRS) [25], Magnetoencephalography (MEG) [26], and Electroencephalography. Each of these methods has its own advantages and disadvantages. For example, fMRI has high spatial resolution [27] but is costly [28] and limited to laboratory environments, while EEG, due to its portability [29], lower cost [30], and high time resolution [5], [9], [31], [32], has become one of the most useful tools for analyzing the cognitive processes and decision-making in consumer neuroscience research [33], [34], [35].
EEG can record the rapid changes in brain activity by measuring the electrical potential difference caused by the activity of neurons [29]. Due to this feature, it is possible to measure neural responses to marketing stimuli and products and detect brain patterns that are related to buying preferences [9]. For this reason, EEG plays an important role in developing consumer behavior prediction models and, when combined with machine learning methods, provides a deeper insight into cognitive and emotional mechanisms [36]. Analyzing EEG signals in this field has usually been done by classical machine learning algorithms such as Support Vector Machine (SVM), Random Forest (RF), and k-Nearest Neighbors (KNN), and deep learning models such as Convolutional Neural Network (CNN) and Long Short-Term Memory (LSTM) [37]. These models usually use features from the time, frequency, and time-frequency domains, and they do not explicitly model the brain’s inter-channel communication structure, while this can be shown as a graph of connections between electrodes, which has a noticeable effect on a better understanding of the decision-making process [38], [39].
@@
This research proposed a framework for consumer behavior prediction by using EEG signals, with the aim of addressing these gaps and introducing several key innovations as follows:
@@
The research is structured as follows: In the Related Works section, previous studies and gaps will be explained. In the Methodology, the dataset, EEG preprocessing, and models will be introduced. The training settings and models’ architectures will be explained in the Experiments section. In the Results section, the results of evaluating models with different metrics will be discussed, and a comparison between the models will be provided. The Conclusions section will focus on the findings, limitations, and future directions of this study.
@@ 2 Related works
In recent decades, neuromarketing research has increasingly utilized novel methods to understand the consumer decision-making process. EEG is recognized as a key tool in the field, due to its high temporal resolution, cost-effectiveness, and lower complexity compared to imaging methods. Previous studies tried to use time domain, frequency domain, or time-frequency domain features of EEG signals to extract indicators related to attention, motivation, or product evaluation and correlate them with behavioral variables.
@@ 2.3 Models: Classical, Ensemble, and Neural Networks
2.4 Research gaps and innovations of this research
@@ 3 Methodology
Figure 1: shows the summary of the introduced method
@@ 3.2.3 Classical Models Descriptions
The diversity of these models makes it possible to analyze the performance of different algorithmic categories (linear, tree-based, distance-based, and probabilistic). Each model was trained independently on different dimensionality reduction pipelines to analyze the combined effects of different pipelines and models.
@@ 3.2.4 Stacking Ensemble Models
To combine the strengths of different models, a stacking-based ensemble model was used. In this method, four different models were trained, and their outputs were fed into a meta-level model. The base models included LR, KNN, NB, and LightGBM to maximize structural diversity. In the final stage, a powerful gradient-boosting model, XGBoost, was employed as the meta-model to make the final prediction.
This structure could consider complex nonlinear interactions and non-local relationships in the data, in addition to simple patterns. In this way, the stacking ensemble method can play an important role in improving the model’s performance.
@@ 3.3.1 Feature Extraction
For the neural network-based method, at first spectrum features of EEG signals were extracted, including detailed spectral indices in the delta, theta, alpha, beta, and gamma bands. These features were computed from both the Fourier domain (FFT) and the power spectral density (PSD). Similar to the classical models’ feature extraction method, a third-order Butterworth bandpass filter and a channel segmentation method were implemented.
@@ 3.3.2 Brain Connectivity Modelling
To model the functional connections between different brain regions, each electrode was represented as a graph node, and edges were calculated by using the Pearson correlation coefficient. This coefficient is defined as follows:
@@
The result of these calculations was a correlation matrix representing the connectivity between every pair of electrodes. This graph-based representation not only considers spectral features but also models the interactions and connections between regions.
@@ 3.3.3 GNN Architectures and Model Descriptions
After constructing the brain graphs, several graph neural network architectures were utilized to model the graphs and analyze them. These models contain baseline models and layers such as Graph Convolutional Network (GCN), which can generate new representations for nodes by applying smoothing operations over their neighbourhoods. This architecture can extract the local dependencies in the graph structure and provide a representation of the entire graph. GraphSAGE employed a sampling and aggregation strategy that is suitable for large-scale data. Graph Attention Networks (GAT) used an attention mechanism to dynamically weight the importance of edges, and allowed the model to detect the significance of the role of different brain regions.
In addition to baseline models, more advanced architectures were also applied. A Residual architecture was used with the skip connections mechanism to facilitate gradient flow and prevent information degradation in deeper networks. Furthermore, multilayer models were proposed, which combine different types of layers to take advantage of different architectures and their combination. In addition, lighter architectures were used to reduce computational complexity, and deeper architectures were utilized to enhance the representational capacity of the models.
@@ 4.3.3 Training Settings
Table 1: Summary of evaluated GNN architectures
@@ 5.1.1 Pipeline A
In the PCA-based pipeline, tree-based and kernel-based methods (such as Random Forest and SVM with RBF kernel) achieved the highest overall accuracy (the accuracy of Random Forest was reported to be around 0.79, and SVM-RBF was around 0.78). However, examination of the discriminant metrics showed that such performance was largely due to the high accuracy in identifying the major class (class 0). For example, Random Forest was able to correctly identify almost all class 0 samples (Recall $\approx 0.99$), but practically failed to identify the minority class (class 1) (Recall $\approx 0.01$). This pattern indicates a strong bias of the models towards the major class and their inefficiency in predicting purchases (minority class). Furthermore, Logistic Regression and Naïve Bayes, in contrast, showed a more balanced performance between the two classes (Class 1 Recall around 0.44), although their overall accuracy was lower. These results indicate that in unbalanced data, the accuracy measure alone cannot reflect the true power of the models in identifying important but small classes. Table 2 provides the summary of PCA-based pipeline results.
Table 2: Summary of PCA-based pipeline results
@@ 5.1.2 Pipeline B
Dimensionality reduction with UMAP had different effects than PCA. In some models, such as Logistic Regression, the use of UMAP resulted in a relatively improved sensitivity to the minority class (Class 1 Recall of about 0.51 compared to 0.44 in PCA). This suggests that UMAP may have been able to reveal some of the nonlinear structures in the data that are useful for detecting purchases. However, the overall pattern remained the same; many models, such as Gaussian Process and SVM-RBF, had high overall accuracy (about 0.79) but very low sensitivity to the minority (Class 1 Recall $\approx 0.02$). Therefore, it can be concluded that although UMAP helps to improve minority detection in some cases, it alone cannot solve the problem of imbalanced data, and the choice of model and optimization criteria plays a more decisive role. In Table 3 the results of the UMAP-based method can be seen.
Table 3: Summary of UMAP-based pipeline results
@@ 5.1.3 Pipeline C
The use of statistical feature selection (t-test) and then dimensionality reduction with PCA led to improvements in Logistic Regression, with a better overall accuracy and a better performance than the PCA-based and similar to the UMAP-based method on minority class detection. Logistic Regression in this case recorded a class 1 recall of 0.51, while its overall accuracy remained around 0.64. This finding suggests that noise removal and focusing on discriminant features can increase the sensitivity of the model to minorities. However, Random Forest still provided higher overall accuracy (Accuracy $\approx 0.78$), but their minority class recall remained very low. These results suggest that feature selection alone is not able to eliminate bias towards the dominant class, although it can be effective for models that are more sensitive to data structure (such as logistic regression). The summary of top features-based method is provided in the Table 4.
Table 4: Summary of top features-based pipeline results
@@ 5.2 Results of Ensemble Models
In general, it can be concluded that the combined pipelines have been able to enhance the performance of the model by increasing the resolution of the data, especially at the level of features related to the major class. Although these approaches have had a slight effect on improving the minority class-related metrics, there is still a significant gap with a balanced classification. Therefore, the results obtained show that although the use of dimensionality reduction and feature selection can change the performance of hybrid models, complementary methods such as resampling or data imbalance-sensitive algorithms are needed to achieve optimal performance in unbalanced data. The results of the ensemble models are available in Table 5.
Table 5: Summary of ensemble models results
@@ 5.3 Results of Graph Neural Network Models
The results obtained from the GNNs (Table 6) show that these approaches have been able to provide acceptable performance in extracting structural features from EEG data, but still face challenges in minority class classification. As can be seen, the overall accuracy of the models is mainly in the range of 0.62 to 0.69. Among them, ResidualGCN with an accuracy of about 0.69 and DeepGNN with an accuracy of about 0.67 have performed better than other models, indicating that increasing the network depth or using residual connections can have a positive role in improving generalizability. In contrast, baseline models such as BaselineGCN, BaselineGAT, and BaselineSAGE have obtained lower accuracy, which indicates their limitations in learning complex relationships between nodes.
@@
Table 6: Summary of graph neural network models results

### v2

@@ 1 Introduction
As purchasing decisions and consumer behaviour play a crucial role and have a significant impact on the success or failure of a business, a deeper understanding of this process has become particularly important for marketing [13], [14], [15], [16] and product development [17], [18]. It is necessary for companies to gain knowledge about their consumers behavior to optimize their marketing strategies [19] and develop and better user experience [20] for costumers, especially in today’s competitive markets [21]. Furthermore, by gaining insight of consumers opinion about new products or marketing campaigns, companies can make competitive advantages and reduce their costs [22].
Many different tools and method have been utilized to measure brain activity or brain response to a stimulus in tasks related consumer neuroscience. These methods include Functional magnetic resonance imaging (fMRI) [23], [24], Functional near-infrared spectroscopy (fNIRS) [25], Magnetoencephalography (MEG) [26], and Electroencephalography. Each of these methods has its own advantages and disadvantages. For example, fMRI has high spatial resolution [27] but it is expensive [28] and limited to laboratory environments, while EEG, due to its portability [29], lower cost [30], and high time resolution [5], [9], [31], [32], has become one of the most useful tools for measuring the brain activity related to decision-making in consumer neuroscience [33], [34], [35].
EEG can record the rapid changes in brain activity by measuring the electrical potential difference caused by the activity of neurons [29]. Due to this feature, it is possible to measure the brain responses to marketing stimuli and products, and detect brain patterns that are related to buying preferences [9]. Because of this reason, EEG plays an important role in developing consumer behavior prediction and, the combination with machine learning methods, can provides a deeper insight into the mechanisms which are related to the cognitive process [36].
Analyzing EEG signals in this field has usually been done by classical machine learning algorithms such as Support Vector Machine (SVM), Random Forest (RF), and k-Nearest Neighbors (KNN), and deep learning models such as Convolutional Neural Network (CNN) and Long Short-Term Memory (LSTM) [37]. These models usually use features from the time, frequency, and time-frequency domains, and they do are not focusing on modelling the interactions between brain channels, while these interactions can be shown as a graph in which node are electrode and edges are the interactions. The graph-based modelling can be helpful for a better understanding of the decision-making process [38], [39].
@@
This research proposed a framework for consumer behavior prediction by using EEG signals, with the aim of addressing these gaps and introducing several key contributions as follows:
@@
The research is structured as follows: In the Related Works section, previous studies and gaps will be explained. In the Methodology, the dataset, EEG preprocessing, and models will be introduced. The training settings and models’ architectures will be explained in the Experiments section. In the Results section, the results of evaluating models with different metrics will be discussed, and a comparison between the models will be provided. The Discussion section will provide theoretical and marketing implications, and Conclusions section will focus on the findings, limitations, and future directions of this study.
@@ 2 Related works
In recent decades, neuromarketing research has increasingly utilized novel methods to understand the consumer decision-making process. EEG is recognized as a key tool in the field, due to its high temporal resolution, cost-effectiveness, and lower complexity compared to imaging methods. Previous studies tried to use time domain, frequency domain, or time-frequency domain features of EEG signals to extract indicators related to attention, motivation, or product evaluation and correlate them with behavioral variables. A summary of the previous key studies is provided in Table  1.
@@ 2.3 Models: Classical, Ensemble, and Neural Networks
Table 1: Summary of EEG-based consumer behavior prediction studies
Ref. No. Year Model Dataset Key Findings
[51] 2021 Artificial Neural Network (ANN), SVM EEG signals from 11 subjects, viewing 20 smartphone images The combination of PCA, Artificial Bee Colony, and ANN achieved the highest accuracy with an average 72%.
[52] 2022 KNN, SVM EEG dataset collected from 15 subjects, viewing 25 sports shoes Differential Entropy features with KNN achieved the highest accuracy (94.22%).
[54] 2022 SVM, KNN, Naive Bayes (NB), Decision Tree (DT) EEG signals from 20 subjects, viewing 3 types of advertisements SVM achieved the highest accuracy with 87% for predicting affective attitude and 84% for predicting purchase intention.
[63] 2023 Deep Learning Network (spectral, spatial, temporal layers) EEG data from 213 subjects, viewing 72 product images Deep learning model achieved about 85% prediction accuracy.
[71] 2023 KNN, SVM, Bidirectional-LSTM EEG data from 25 subjects, viewing 42 product images Bidirectional-LSTM achieved the highest accuracy (96.83%) using frontal region signals.
[75] 2024 CNN–LSTM, SVM, RF, CNN-Transformer EEG dataset collected from 25 subjects, viewing 42 product images CNN–LSTM model achieved the highest accuracy (97.12%).
[59] 2024 Gradient Boost (GB), NB, KNN, DT, SVM, RF, CNN EEG data from 33 participants, exposed to scented and unscented product boxes SVM with Lempel–Ziv Complexity feature achieved 93% accuracy, while CNN achieved 91.6% accuracy.
[53] 2024 SVM, RF, KNN, Shallow Neural Network EEG data from 66 participants, for 5 products across 6 categories SVM achieved the highest accuracy (87.1%) in predicting purchasing decisions.
[…]
2.4 Research gaps and contributions of this research
@@ 3 Methodology
Figure 1: Overview of the proposed methodology for EEG-based consumer behavior prediction
@@ 3.2.3 Classical Models Descriptions
The diversity of these models makes it possible to analyze the performance of different algorithmic categories (linear, tree-based, distance-based, and probabilistic). The models were trained using different pipelines with different dimensionality reduction and feature selection methods to provide an analysis of the models’ performance in each pipeline.
@@ 3.2.4 Stacking Ensemble Models
To benefit from the strengths of different classical models, a stacking ensemble model was utilized to combine the models’ predictions. Four different models were trained, and the final predictions were made by a meta model. Each of the four base model make an initial prediction, and the meta model predicts by feeding from the base models’ predictions. Diverse base models have been used, including LR, KNN, NB, and LightGBM, which increases the variety of prediction methods. Furthermore, XGBoost has been utilized as the meta model to provide the final predictions.
By using this architecture for the stacking model, it can be able to model complex non-linear patterns, as well as the linear interactions. These stacking models can provide a deeper insight into combining different models to improve performance.
@@ 3.3.1 Feature Extraction
In contrast with the classical model features, the spectral features of EEG signals were extracted and used for training the GNNs. These include the Fourier domain (FFT) and the power spectral density (PSD) features from the delta, theta, alpha, beta, and gamma bands. Similar to the classical models’ feature extraction method, a third-order Butterworth bandpass filter and a channel segmentation method were implemented.
@@ 3.3.2 Brain Connectivity Modelling
For modelling the brain connectivity and constructing graphs, the EEG electrodes were considered as graph nodes, and the Pearson correlation coefficient was used to calculate the edge features. This coefficient is defined as follows:
@@
As a result of this, a correlation matrix has been constructed for each data set, which represents the connectivity between each pair of nodes. This can help consider both spectral features and the interaction between brain regions.
@@ 3.3.3 GNN Architectures and Model Descriptions
A wide range of graph neural network architectures were implemented to analyze the graphs which have been constructed using EEG-brain features and Pearson correlation coefficient. These models contain baseline models and layers such as Graph Convolutional Network (GCN), Graph Attention Networks (GAT) and GraphSAGE. While GCN can represent new features for each node by using the local connections and nearby nodes, GAT uses an attention mechanism to put more weight on important edges which can define the importance of each brain region. Also, GraphSAGE uses a method that have a significant impact on graphs with high dimension node features.
In addition to baseline models, more advanced architectures were also utilized for classifying the consumer decision. A Residual architecture with the skip connections mechanism was used to reduce the information loss in deeper layers. Also, multilayer models were introduced, which combine different types of layers to take advantage of different architectures and their combination. Furthermore, lighter architectures were used to reduce computational complexity, and deeper architectures were utilized to enhance the representational capacity of the models.
@@ 4.3.3 Training Settings
Table 2: Description of the GNN architectures used in this study
@@ 5.1.1 Pipeline A
In the PCA-based pipeline, tree-based and kernel-based methods (such as Random Forest and SVM with RBF kernel) achieved the highest overall accuracy (the accuracy of Random Forest was reported to be around 0.79, and SVM-RBF was around 0.78). However, examination of the discriminant metrics showed that such performance was largely due to the high accuracy in identifying the major class (class 0). For example, Random Forest was able to correctly identify almost all class 0 samples (Recall $\approx 0.99$), but practically failed to identify the minority class (class 1) (Recall $\approx 0.01$). This pattern indicates a strong bias of the models towards the major class and their inefficiency in predicting purchases (minority class). Furthermore, Logistic Regression and Naïve Bayes, in contrast, showed a more balanced performance between the two classes (Class 1 Recall around 0.44), although their overall accuracy was lower. These results indicate that in unbalanced data, the accuracy measure alone cannot reflect the true power of the models in identifying important but small classes. Figure 2 and Table 3 provide a summary of PCA-based pipeline results.
Figure 2: Comparison of classical machine learning models with the PCA-based pipeline
Table 3: Performance of the PCA-based machine learning pipeline across different models
@@ 5.1.2 Pipeline B
Dimensionality reduction with UMAP had different effects than PCA. In some models, such as Logistic Regression, the use of UMAP resulted in a relatively improved sensitivity to the minority class (Class 1 Recall of about 0.51 compared to 0.44 in PCA). This suggests that UMAP may have been able to reveal some of the nonlinear structures in the data that are useful for detecting purchases. However, the overall pattern remained the same; many models, such as Gaussian Process and SVM-RBF, had high overall accuracy (about 0.79) but very low sensitivity to the minority (Class 1 Recall $\approx 0.02$). Therefore, it can be concluded that although UMAP helps to improve minority detection in some cases, it alone cannot solve the problem of imbalanced data, and the choice of model and optimization criteria plays a more decisive role. In Figure 3 and Table 4 the results of the UMAP-based method can be seen.
Figure 3: Comparison of classical machine learning models with the UMAP-based pipeline
Table 4: Performance of the UMAP-based machine learning pipeline across different models
@@ 5.1.3 Pipeline C
The use of statistical feature selection (t-test) and then dimensionality reduction with PCA led to improvements in Logistic Regression, with a better overall accuracy and a better performance than the PCA-based and similar to the UMAP-based method on minority class detection. Logistic Regression in this case recorded a class 1 recall of 0.51, while its overall accuracy remained around 0.64. This finding suggests that noise removal and focusing on discriminant features can increase the sensitivity of the model to minorities. However, Random Forest still provided higher overall accuracy (Accuracy $\approx 0.78$), but their minority class recall remained very low. These results suggest that feature selection alone is not able to eliminate bias towards the dominant class, although it can be effective for models that are more sensitive to data structure (such as logistic regression). The summary of top features-based method is provided in Figure 4 and Table 5.
Figure 4: Comparison of classical machine learning models with the top features-based pipeline
Table 5: Performance of the top features-based machine learning pipeline across different models
@@ 5.2 Results of Ensemble Models
In general, it can be concluded that the combined pipelines have been able to enhance the performance of the model by increasing the resolution of the data, especially at the level of features related to the major class. Although these approaches have had a slight effect on improving the minority class-related metrics, there is still a significant gap with a balanced classification. Therefore, the results obtained show that although the use of dimensionality reduction and feature selection can change the performance of ensemble models, complementary methods such as resampling or data imbalance-sensitive algorithms are needed to achieve optimal performance in unbalanced data. The results of the ensemble models are available in Figure 5 and Table 6.
Figure 5: Comparison of ensemble models with PCA-based, UMAP-based, and top features-based pipelines
Table 6: Performance of the ensemble machine learning models across different pipelines
@@ 5.3 Results of Graph Neural Network Models
The results obtained from the GNNs (Figure 6 and Table 7) show that these approaches have been able to provide acceptable performance in extracting structural features from EEG data, but still face challenges in minority class classification. As can be seen, the overall accuracy of the models is mainly in the range of 0.62 to 0.69. Among them, ResidualGCN with an accuracy of about 0.69 and DeepGNN with an accuracy of about 0.67 have performed better than other models, indicating that increasing the network depth or using residual connections can have a positive role in improving generalizability. In contrast, baseline models such as BaselineGCN, BaselineGAT, and BaselineSAGE have obtained lower accuracy, which indicates their limitations in learning complex relationships between nodes.
@@
Figure 6: Comparison of GNN models with brain connectivity features
Table 7: Performance of the GNN machine learning models
@@
5.4 Comparative Analysis of Modelling Approaches
The results obtained show that the three main categories of approaches used, namely classical models based on EEG statistical features, ensemble models with feature selection or dimensionality reduction, and models based on graph neural networks (GNNs) using spectral features, each have their own strengths and weaknesses. Classical models were able to provide relatively stable classification performance using statistical features extracted from EEG signals and achieved noticeable results in major class detection. These models are computationally lighter and, due to the simplicity of their structure, have relatively good generalization capabilities in limited data conditions, but they have limitations in learning complex and nonlinear patterns between EEG channels.
Ensemble models can show a balance in terms of accuracy and minority class prediction, while they are using different dimensionality reduction methods, such as PCA and UMAP, proposed in the pipelines. The stacking model, which used UMAP as a dimensionality reduction method, achieved higher accuracy than other ensemble models that were trained in the different pipelines. However, like classical models, these models still faced challenges in identifying the minority class.
Unlike the classical and ensemble models, GNN-based models, which used EEG spectral features and constructed brain connectivity graphs for training, were able to provide deeper insights in terms of the interactions between the nodes. However, their overall accuracy remained lower than the ensemble models in most cases, and only some more advanced architectures (such as ResidualGCN and DeepGNN) were able to provide more competitive results. Overall, the minority class prediction results in graph-based models are more remarkable than other methods, and in cases such as BalancedGAT, it was able to obtain higher Recall for the minority class.
Overall, classical and ensemble models based on EEG statistical features still succeeded in providing better accuracy, while GNN models, despite lower accuracy, have higher potential in learning complex relationships between channels and exploiting spectral structures, and providing a better performance on minor class prediction. The results suggest that in future research combination of classical approaches with graph-based architectures can be utilized to develop hybrid methods, which can take the advantages of both methods and improve the predictions for the minority class.
6 Discussion
The obtained results from the utilized models show that classical, ensemble, and GNN models each have their own abilities in analysing and modelling EEG signals collected by marketing stimuli. Overall, classical models show a better accuracy, while GNNs show a better performance in identifying the minority class by modelling the structural relationships between brain channels. This research shows that GNN models performed better in identifying complex cognitive patterns, furthermore, their performance on the minority class is better than the classical and ensemble models that combined with different feature selection models.
6.1 Theoretical Implications
From a theoretical perspective, this study shows that graph-based modelling can be introduced as a novel method in neuromarketing tasks. Instead of focusing on time domain and frequency domain features, using brain connectivity features can provide a deeper insight into the consumer decision-making process. This can be utilized in the development of data-driven cognitive models in marketing and neuroscience.
GNNs are able to model the topological connections between electrodes that collect the EEG data from the brain. Due to this characteristic of GNNs, they can model the interactions between different nodes of the graph, which are the electrodes in an EEG-based task. These research findings introduce a new idea in neuromarketing and give a direction to future neuromarketing studies.
Overall, by introducing a graph-based framework for EEG analysis in neuromarketing, this study provides a new approach for modelling consumer brain data collected by using marketing stimuli. These findings can lead to developing novel theoretical models that can explain the consumer decision-making process as a brain connectivity network.
6.2 Practical Implications
From a practical perspective, the results of this study can provide a deeper insight for companies, brands, and marketing specialists. The proposed models have the ability to classify consumers’ brain EEG signals in response to a marketing stimulus. This can lead to new developments for optimizing marketing campaigns and product design.
In addition, using machine learning models could be helpful for companies to segment their consumers into different groups by using brain responses. Thus, personalized marketing can be implemented on a deeper level. Furthermore, companies can use EEG-based methods to assess the effectiveness of pre-launch advertising by showing multiple versions of an advertisement to different groups and using EEG to select the most effective version.